# ACT-ing on inflation: Implications of non Bunch-Davies initial condition and reheating on single-field slow roll models

### v1

@@ I Introduction
The Planck 2018 data constrained the scalar spectral index to $n_{{}_{\mathrm{S}}}=0.9651\pm 0.0044$ [5, 6]. An improved constraint, combining Planck data with measurements from the South Pole Telescope (SPT), yielded $n_{{}_{\mathrm{S}}}=0.9647\pm 0.0037$ [7]. Recently, the latest data released by the Atacama Cosmology Telescope (ACT) [8, 9], when combined with Planck data, has provided a significantly different constraint of $n_{{}_{\mathrm{S}}}=0.9709\pm 0.0038$. Considering the DESI DR1 (P-ACT-LB-BK18), the constraint has further improved to $n_{{}_{\mathrm{S}}}=0.9743\pm 0.0034$ [8]. This value suggests a spectrum that is closer to scale invariance compared to previous results. It is also worth noting that the upper bound on the tensor-to-scalar ratio is $r<0.036$ [10].
The latest constraints on the $n_{{}_{\mathrm{S}}}-r$ plane have renewed interest in re-evaluating existing inflationary models. It has been shown that while the chaotic inflation model with a $\phi^{1/3}$ potential is now consistent with current data, the Starobinsky $R^{2}$ model [11] is nearly pushed outside the $2\sigma$ confidence region [9]. Furthermore, inflation with a $m^{2}\phi^{2}$ potential can also be brought into agreement with observations when a non-minimal coupling to gravity is included [12]. Other theoretically motivated scenarios, such as Higgs inflation [13, 14, 15] and worm inflation [16], and others [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] have also been revisited in light of recent ACT data. Notably, there have been several recent efforts to bring plateau-like models—especially Starobinsky inflation—back into the allowed parameter space, either by incorporating a phase of reheating [25, 27] or by adopting non Bunch-Davies (NBD) initial conditions [30].
@@
The phase of reheating is the intermediate epoch between the end of inflation and the onset of the standard radiation-dominated era, during which the inflaton undergoes a quasi-periodic oscillation and decays into relativistic particles constituting the standard model. The phase of reheating governs the thermal history of the universe and sets the initial conditions for Big Bang Nucleosynthesis (BBN) [44, 45, 46]. Reheating is often characterised by parameters such as the reheating temperature $T_{\mathrm{re}}$, duration, and an effective equation of state (EoS) $w$. As stated, the phase influences the mapping between the predictions from inflation and observables in CMB. Specifically, it affects the number of e-folds between the pivot scale and the end of inflation, thereby shifting the prediction of $n_{{}_{\mathrm{S}}}-r$.
@@
We also consider an additional constraint arising from the overproduction of primordial gravitational waves (GWs) generated by vacuum fluctuations during inflation. The EoS during reheating can modify the shape of the GW modes that re-enter the Hubble radius during this phase [47, 48, 49, 50]. For an EoS with $w<1/3$, the spectrum acquires a red tilt, whereas for $w>1/3$, the tilt becomes blue. Furthermore, $T_{\mathrm{re}}$ and $w$ dictate the duration of reheating, thereby determining how long the modified tilt persists in the GW spectrum. It is important to note that inside the Hubble radius, GWs behave like radiation. An excess of GWs (which arises for $w>1/3$) thus contributes to the effective number of relativistic degrees of freedom, which can be constrained by BBN and CMB [51]. Consequently, these bounds provide a constraint on the duration of reheating or eventually on $T_{\mathrm{re}}$ for a given $w$.
@@
This paper is organized as follows. In Sec. II, we review the framework of slow roll inflation and the generation of primordial perturbations, focusing on the two-point correlations or the power spectrum. We also discuss how a NBD initial state modifies key observables such as the scalar spectral index $n_{{}_{\mathrm{S}}}$ and the tensor-to-scalar ratio $r$. In Sec. III, we examine the post-inflationary phase of reheating, exploring how the reheating parameters affect the duration of inflation and subsequently alter the predictions for $n_{{}_{\mathrm{S}}}$ and $r$. Sec. IV is devoted to the impact of $\Delta{N}_{\mathrm{eff}}$ on the amplitude of primordial gravitational waves. We derive the resulting lower bound on the reheating temperature from current observational constraints. In Sec. V, we analyze a variety of single-field slow roll inflationary models and investigate how the inclusion of the dynamics of reheating and the NBD initial condition influence their compatibility with the latest CMB observations from the P-ACT-LB-BK18 dataset. Finally, in Sec. VI, we present our conclusions along with a brief outlook for future directions.
@@
Let us discuss a few points about the notations. We shall work with natural units such that $\hbar=c=1$, and the reduced Planck mass is $M_{{}_{\mathrm{Pl}}}=\left(8\,\pi\,G\right)^{-1/2}\simeq 2.4\times 10^{18}\,\mathrm{GeV}$. The signature of the metric is followed (,+,+,+)(-,+,+,+)( - , + , + , + ) and the background is spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW). An overdot and an overprime shall denote differentiation with respect to the cosmic time $t$ and the conformal time $\eta$, respectively. Derivatives with respect to any other quantity will be denoted as subscript by the quantity.
@@ II Inflation and primordial scalar perturbations
The parameter $\beta_{k}$ needs to be such that after a certain cutoff scale, the mode restores the BD state. A choice of physical model, which is often used for $\beta_{k}$ is given by [56]
@@
where $M$ is the physical cutoff, which during inflation, needs to be larger than the Hubble parameter. Furthermore, the choice of $\beta$ provides $\mathrm{d}\beta_{k}/\mathrm{d}\ln k=-2\beta_{k}k^{2}/(a^{2}M^{2})$. At this point, it is useful to note that, to stay in the linear regime and to avoid backreaction, $\beta_{k}\ll 1$ is required. In this work, we shall stick to the choice of the cutoff at pivot scale to be $aM\simeq 5k_{\ast}$, and hence, we shall limit choosing $\beta_{0}<0.2$ in order to prevent backreaction. In the next section, we shall discuss the effects due to the phase of reheating on the inflationary observables.
@@ III Constraints from the phase of reheating
As already mentioned, the phase of reheating connects the inflation to the epoch of standard radiation domination. It is a phase with the energy density dominated by inflation, and during which the field transfers its energy through various couplings to the standard model particles. In general, the phase of reheating can be characterized by the following variables: reheating EoS, $w$, reheating temperature, $T_{\mathrm{re}}$, and the duration of reheating, $N_{\mathrm{re}}$. The end of reheating is defined when $\rho_{\phi}=\rho_{{}_{\rm R}}=\rho_{\mathrm{re}}$, where $\rho_{\phi}$ and $\rho_{{}_{\rm R}}$ are the energy density of inflation and radiation respectively. Hence, one can use the conservation of entropy, $g_{\rm s}a^{3}T^{3}=\text{constant}$ at the end of reheating, where $g_{\rm s}$ is the relativistic degrees of freedom associated with the entropy. Using the conservation relation, we get [60, 61]
@@
where we have used, at present day $g_{\rm s}^{0}=43/11$. The quantity $N_{\mathrm{k}}$ is the duration from the pivot scale to the end of inflation, and $H_{{}_{\mathrm{I}}}$ is the Hubble parameter at the end of inflation. Further, in a background driven by a general EoS, $w$, the total energy density falls as $\rho\propto a^{-3(1+w)}$. Hence, using the relation $\rho_{\mathrm{re}}=(\pi^{2}/30)g_{\rm re}T_{\mathrm{re}}^{4}$, the temperature of reheating is given by
$T_{\mathrm{re}}=\left(\frac{90M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}{\pi^{2}g_{\rm re}}\right)^{1/4}\mathrm{e}^{-(3/4)(1+w)N_{\mathrm{re}}}.$ (13)
It is straightforward to see from Eq. (13), that $N_{\mathrm{re}}$, $w$ and $T_{\mathrm{re}}$ are not completely independent. One can be obtained if the other two quantities are specified. Combining Eqs. (12) and (13) we get the duration of inflation in terms of $w$ and $T_{\mathrm{re}}$ to be
$N_{\mathrm{k}}$ $=\log\left[\left(\frac{43}{11\,g_{\rm s}^{\rm re}}\right)^{1/3}\left(\frac{\pi^{2}\,g_{\rm re}}{90M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}\right)^{\frac{1}{3(1+w)}}\right.$
×$\left.\quad\quad\quad\times T_{0}\frac{H_{{}_{\mathrm{I}}}}{k_{\star}}T_{\mathrm{re}}^{\frac{1-3w}{3(1+w)}}\right],$ (14)
where $T_{0}$ is the temperature of the CMB at the present day. From Eq. (14), is straightforward to see that, for $w=1/3$, $N_{\mathrm{k}}$ will be independent of $T_{\mathrm{re}}$. For $w<1/3$, a higher reheating temperature will lead to a higher value of $N_{\mathrm{k}}$, whereas the dependence becomes opposite for $w>1/3$. In slow roll inflation, $N_{\mathrm{k}}$ becomes extremely important to determine the value of $\phi$ at the pivot scale. Which further is used to determine $\epsilon$, $\eta$, and $n_{{}_{\mathrm{S}}}$. The details of obtaining $N_{\mathrm{k}}$ shall be discussed later. In the next section, we shall give a brief overview of the bound of $\Delta{N}_{\mathrm{eff}}$ to the strength of the GW spectrum.
@@ IV Constraints from the primary GWs and $\Delta{N}_{\mathrm{eff}}$
×$\quad\times\Gamma^{2}\left(1+\frac{\nu}{\mu}\right)\left(\frac{k}{2\mu k_{\mathrm{re}}}\right)^{n_{{}_{\mathrm{GW}}}},$ (15)
where the quantities $\mu=(1/2)(1+3w)$ and $\nu=(3/4)(1-w)$. The quantity $c_{g}=(g_{\rm eq}/g_{0})(g_{\rm s}^{\rm eq}/g_{\rm s}^{0})^{4/3}$ contains the combination of relativistic degrees of freedom. Also, $\Omega_{{}_{\mathrm{R}}}$ is the energy density of the radiation today. The spectral index of the GWs is given by $n_{{}_{\mathrm{GW}}}=-2(1-3w)/(1+3w)$. It is clear that during radiation dominated epoch ($w=1/3$), we obtain a scale invariant spectrum. We also see that reheating with $w<1/3$ will induce a red tilt and $w>1/3$ will induce a blue tilt to the spectrum for the modes that enter during the phase of reheating, i.e., for $k_{\mathrm{e}}>k>k_{\mathrm{re}}$.
@@
where $\Omega_{\rm\gamma}h^{2}\simeq 2.47\times 10^{-5}$ is the photon relic density at the present day. It is straightforward to notice that the bound from $\Delta{N}_{\mathrm{eff}}$ becomes effective only for a blue-tilted spectrum which is obtained for a steeper EoS, i.e. $w>1/3$. In such a case, upon substitution of Eq. (15) in Eq. (17) and performing the integration, one can find
$\int_{k_{\mathrm{re}}}^{k_{\mathrm{e}}}\frac{\mathrm{d}k}{k}\Omega_{{}_{\mathrm{GW}}}(k)h^{2}$ $\simeq c_{g}\,\Omega_{{}_{\mathrm{R}}}h^{2}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\Gamma^{2}\left(\frac{5+3w}{2+6w}\right)$
×$\quad\times\frac{(1+3w)^{1+4/(1+3w)}}{2\pi(3w-1)}\left(\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}\right)^{\frac{6w-2}{1+3w}},$
$\simeq c_{g}\,\Omega_{{}_{\mathrm{R}}}h^{2}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\zeta(w)\left(\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}\right)^{\frac{6w-2}{1+3w}}$ (18)
in Eq. (18), the ratio between $k_{\mathrm{e}}$ and $k_{\mathrm{re}}$ can be expressed in terms of $H_{{}_{\mathrm{I}}}$, $w$ and $T_{\mathrm{re}}$ as
$\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}=\left(\frac{90}{\pi^{2}g_{\rm re}}\frac{M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}{T_{\mathrm{re}}^{4}}\right)^{\frac{1+3\,w}{6\,(1+w)}}.$ (19)
@@
×$\quad\times\left(\frac{c_{g}\Omega_{{}_{\mathrm{R}}}h^{2}\,\zeta(w)}{5.61\times 10^{-6}\,\Delta N_{\rm eff}}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\right)^{\frac{3\,(1+w)}{4\,(3\,w-1)}}.$ (20)
In Eq. (20), $\Delta{N}_{\mathrm{eff}}$ is tightly constrained by recent ACT data, yielding an upper limit of $\Delta{N}_{\mathrm{eff}}\leq 0.17$ at $95\%$ confidence level [8, 9]. It is important to note that there also exists a lower bound on the reheating temperature, set by BBN, which requires $T_{\mathrm{re}}\gtrsim 4$ MeV. It can be shown that the bound on $T_{\mathrm{re}}$ derived from Eq. (20) is stronger than the BBN bound only for equations of state with $w>0.58$. Therefore, in our analysis, we consider the minimum reheating temperature to be max[$T_{\mathrm{re}},T_{{}_{\mathrm{BBN}}}$], while the maximum reheating temperature corresponds to the case of instantaneous reheating, which can be directly computed from the energy scale at the end of inflation. With these tools in hand, we now turn to a discussion of the slow roll models of interest in the next section.

### v2

@@ I Introduction
[…]
In [17], the authors varied, in addition to the six standard cosmological parameters, extra parameters such as the running of the scalar spectral index, the tensor-to-scalar ratio, and the spatial curvature. They demonstrated that these additional parameters have only a weak impact on the upper bound of $r$, whereas $n_{{}_{\mathrm{S}}}$ is significantly affected. In particular, a value of $n_{{}_{\mathrm{S}}}\sim 0.97$ can be obtained using the Planck18 + BK15 data, and an even higher value of $n_{{}_{\mathrm{S}}}$ is achievable when WMAP data are combined with ACTPol and SPT3G. The authors of [18] demonstrated that the Starobinsky, Higgs, and $\alpha$-attractor models provide a good fit to the Planck18 + BK18 data. [19] discusses that, with suitable reheating parameters, $n_{{}_{\mathrm{S}}}$ can reach $\sim 0.97$ with $n\to\infty$ in the $\alpha$-attractor model.
The latest constraints (after ACT) on the $n_{{}_{\mathrm{S}}}-r$ plane have renewed interest in re-evaluating existing inflationary models. It has been shown that while the chaotic inflation model with a $\phi^{1/3}$ potential is now consistent with current data, the Starobinsky $R^{2}$ model [20] is nearly pushed outside the $2\sigma$ confidence region [9]. Furthermore, inflation with a $m^{2}\phi^{2}$ potential can also be brought into agreement with observations when a non-minimal coupling to gravity is included [21]. Other theoretically motivated scenarios, such as Higgs inflation [22, 23, 24, 25, 26] and warm inflation [27], and others [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] have also been revisited in light of recent ACT data. Notably, there have been several recent efforts to bring plateau-like models—especially Starobinsky inflation—back into the allowed parameter space, either by incorporating a phase of reheating [36, 38] or by adopting non Bunch-Davies (NBD) initial conditions [42].
@@
The phase of reheating is the intermediate epoch between the end of inflation and the onset of the standard radiation-dominated era, during which the inflaton undergoes a quasi-periodic oscillation and decays into relativistic particles constituting the standard model (SM). The phase of reheating governs the thermal history of the universe and sets the initial conditions for Big Bang Nucleosynthesis (BBN) [56, 57, 58]. Reheating is often characterised by parameters such as the reheating temperature $T_{\mathrm{re}}$, duration, and an effective equation of state (EoS) $w_{\mathrm{re}}$. As stated, the phase influences the mapping between the predictions from inflation and observables in CMB. Specifically, it affects the number of e-folds between the pivot scale and the end of inflation, thereby shifting the prediction of $n_{{}_{\mathrm{S}}}-r$.
@@
We also consider an additional constraint arising from the overproduction of primordial gravitational waves (GWs) generated by vacuum fluctuations during inflation. The EoS during reheating can modify the shape of the GW modes that re-enter the Hubble radius during this phase [59, 60, 61, 62]. For an EoS with $w_{\mathrm{re}}<1/3$, the spectrum acquires a red tilt, whereas for $w_{\mathrm{re}}>1/3$, the tilt becomes blue. Furthermore, $T_{\mathrm{re}}$ and $w_{\mathrm{re}}$ dictate the duration of reheating, thereby determining how long the modified tilt persists in the GW spectrum. It is important to note that inside the Hubble radius, GWs behave like radiation. An excess of GWs (which arises for $w_{\mathrm{re}}>1/3$) thus contributes to the effective number of relativistic degrees of freedom, which can be constrained by BBN and CMB [63]. Consequently, these bounds provide a constraint on the duration of reheating or eventually on $T_{\mathrm{re}}$ for a given $w_{\mathrm{re}}$.
@@
This paper is organized as follows. In Sec. II, we review the framework of slow roll inflation and the generation of primordial perturbations, focusing on the two-point correlations or the power spectrum. We also discuss how a NBD initial state modifies key observables such as the scalar spectral index $n_{{}_{\mathrm{S}}}$ and the tensor-to-scalar ratio $r$. In Sec. III, we examine the post-inflationary phase of reheating, exploring how the reheating parameters affect the duration of inflation and subsequently alter the predictions for $n_{{}_{\mathrm{S}}}$ and $r$. Sec. IV is devoted to the impact of $\Delta{N}_{\mathrm{eff}}$ on the amplitude of primordial gravitational waves. We derive the resulting lower bound on the reheating temperature from current observational constraints. In Sec. V, we analyze a variety of single-field slow roll inflationary models and investigate how the inclusion of the dynamics of reheating and the NBD initial condition influence their compatibility with the latest CMB observations from the P-ACT-LB-BK18 dataset. Finally, in Sec. VII, we present our conclusions along with a brief outlook for future directions.
@@
Let us discuss a few points about the notations. We shall work with natural units such that $\hbar=c=1$, and the reduced Planck mass is $M_{{}_{\mathrm{Pl}}}=\left(8\,\pi\,G\right)^{-1/2}\simeq 2.4\times 10^{18}\,\mathrm{GeV}$. The signature of the metric is followed (,+,+,+)(-,+,+,+) and the background is spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW). An overdot and an overprime shall denote differentiation with respect to the cosmic time $t$ and the conformal time $\eta$, respectively. Derivatives with respect to any other quantity will be denoted as subscript by the quantity.
@@ II Inflation and primordial scalar perturbations
The parameter $\beta_{k}$ needs to be such that after a certain cutoff scale, the mode restores the BD state. A choice of physical model, which is often used for $\beta_{k}$ is given by [68, 72, 73, 42]
@@
where $M$ is the physical cutoff, which during inflation, needs to be larger than the Hubble parameter. Furthermore, the choice of $\beta_{k}$ provides $\mathrm{d}\beta_{k}/\mathrm{d}\ln k=-2\beta_{k}k^{2}/(a^{2}M^{2})$ [72, 73, 42]. At this point, it is useful to note that, to stay in the linear regime and to avoid backreaction, $\beta_{k}\ll 1$ is required [42]. In [74, 75, 76], the author demonstrated that the back reaction is not very constraining for corrections to the power spectrum as long as the scale is different from the Planck scale. The upper value on $\beta_{k}$ is estimated to be $HM_{{}_{\mathrm{Pl}}}\sqrt{\epsilon\eta^{\prime}}/M^{2}$. Hence, a large value of $\beta_{0}$ can be obtained by changing the value of $H/M$. In this work, we shall stick to the choice of the cutoff at pivot scale to be $aM\simeq 5k_{\ast}$ to ensure that the pivot scale is well inside the Hubble radius at the initial time [42]. Further, we shall limit choosing $\beta_{0}<0.2$ to avoid backreaction (a discussion regarding the choice of the value is given in Sec. VI). In the next section, we shall discuss the effects due to the phase of reheating on the inflationary observables.
@@ III Constraints from the phase of reheating
As already mentioned, the phase of reheating connects the inflation to the epoch of standard radiation domination. It is a phase with the energy density dominated by inflation, and during which the field transfers its energy through various couplings to the SM particles. There exist several processes through which the inflaton can transfer its energy to daughter particles. This transfer can occur either via non-gravitational couplings [77], or through purely gravitational interactions [78, 79, 80, 19] (where no additional coupling is required and reheating occurs purely due to time-dependent and curved background). In addition, ultralight primordial black holes may form during reheating, which can subsequently emit SM particles through Hawking evaporation [81, 62]. In this work, we focus on studying the reheating phase in a model-independent manner, without delving into the specifics of the couplings. In general, the phase of reheating can be characterized by the following variables: reheating EoS, $w_{\mathrm{re}}$, reheating temperature, $T_{\mathrm{re}}$, and the duration of reheating, $N_{\mathrm{re}}$. The end of reheating is defined when $\rho_{\phi}=\rho_{{}_{\rm R}}=\rho_{\mathrm{re}}$, where $\rho_{\phi}$ and $\rho_{{}_{\rm R}}$ are the energy density of inflation and radiation respectively. Hence, one can use the conservation of entropy, $g_{\rm s}a^{3}T^{3}=\text{constant}$ at the end of reheating, where $g_{\rm s}$ is the relativistic degrees of freedom associated with the entropy. Using the conservation relation, we get [82, 83]
@@
where we have used, at present day $g_{\rm s}^{0}=43/11$. The quantity $N_{\mathrm{k}}$ is the duration from the pivot scale to the end of inflation, and $H_{{}_{\mathrm{I}}}$ is the Hubble parameter at the end of inflation. Further, in a background driven by a general EoS, $w_{\mathrm{re}}$, the total energy density falls as $\rho\propto a^{-3(1+w_{\mathrm{re}})}$. Hence, using the relation $\rho_{\mathrm{re}}=(\pi^{2}/30)g_{\rm re}T_{\mathrm{re}}^{4}$, the temperature of reheating is given by
$T_{\mathrm{re}}=\left(\frac{90M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}{\pi^{2}g_{\rm re}}\right)^{1/4}\mathrm{e}^{-(3/4)(1+w_{\mathrm{re}})N_{\mathrm{re}}}.$ (13)
It is straightforward to see from Eq. (13), that $N_{\mathrm{re}}$, $w_{\mathrm{re}}$ and $T_{\mathrm{re}}$ are not completely independent. One can be obtained if the other two quantities are specified. Combining Eqs. (12) and (13) we get the duration of inflation in terms of $w_{\mathrm{re}}$ and $T_{\mathrm{re}}$ to be
$N_{\mathrm{k}}$ $=\log\left[\left(\frac{43}{11\,g_{\rm s}^{\rm re}}\right)^{1/3}\left(\frac{\pi^{2}\,g_{\rm re}}{90M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}\right)^{\frac{1}{3(1+w_{\mathrm{re}})}}\right$.
$\left.\quad\quad\quad\times T_{0}\frac{H_{{}_{\mathrm{I}}}}{k_{\star}}T_{\mathrm{re}}^{\frac{1-3w_{\mathrm{re}}}{3(1+w_{\mathrm{re}})}}\right],$ (14)
where $T_{0}$ is the temperature of the CMB at the present day. From Eq. (14), is straightforward to see that, for $w_{\mathrm{re}}=1/3$, $N_{\mathrm{k}}$ will be independent of $T_{\mathrm{re}}$. For $w_{\mathrm{re}}<1/3$, a higher reheating temperature will lead to a higher value of $N_{\mathrm{k}}$, whereas the dependence becomes opposite for $w_{\mathrm{re}}>1/3$. In slow roll inflation, $N_{\mathrm{k}}$ becomes extremely important to determine the value of $\phi$ at the pivot scale. Which further is used to determine $\epsilon$, $\eta$, and $n_{{}_{\mathrm{S}}}$. The details of obtaining $N_{\mathrm{k}}$ shall be discussed later. In the next section, we shall give a brief overview of the bound of $\Delta{N}_{\mathrm{eff}}$ to the strength of the GW spectrum.
@@ IV Constraints from the primary GWs and $\Delta{N}_{\mathrm{eff}}$
$\quad\times\Gamma^{2}\left(1+\frac{\nu}{\mu}\right)\left(\frac{k}{2\mu k_{\mathrm{re}}}\right)^{n_{{}_{\mathrm{GW}}}},$ (15)
where the quantities $\mu=(1/2)(1+3w_{\mathrm{re}})$ and $\nu=(3/4)(1-w_{\mathrm{re}})$. The quantity $c_{g}=(g_{\rm eq}/g_{0})(g_{\rm s}^{\rm eq}/g_{\rm s}^{0})^{4/3}$ contains the combination of relativistic degrees of freedom. Also, $\Omega_{{}_{\mathrm{R}}}$ is the energy density of the radiation today. The spectral index of the GWs is given by $n_{{}_{\mathrm{GW}}}=-2(1-3w_{\mathrm{re}})/(1+3w_{\mathrm{re}})$. It is clear that during radiation dominated epoch ($w_{\mathrm{re}}=1/3$), we obtain a scale-invariant spectrum. We also see that reheating with $w_{\mathrm{re}}<1/3$ will induce a red tilt and $w_{\mathrm{re}}>1/3$ will induce a blue tilt to the spectrum for the modes that enter during the phase of reheating, i.e., for $k_{\mathrm{e}}>k>k_{\mathrm{re}}$. At this stage, we clarify that the exact form of the tensor power spectrum can be obtained by considering the evolution of a specific model [85, 86]. Here we have considered the de Sitter approximation, under which inflation is assumed to generate a scale-invariant tensor power spectrum throughout.
@@
where $\Omega_{\rm\gamma}h^{2}\simeq 2.47\times 10^{-5}$ is the photon relic density at the present day. It is straightforward to notice that the bound from $\Delta{N}_{\mathrm{eff}}$ becomes effective only for a blue-tilted spectrum which is obtained for a steeper EoS, i.e. $w_{\mathrm{re}}>1/3$. In such a case, upon substitution of Eq. (15) in Eq. (17) and performing the integration, one can find
$\int_{k_{\mathrm{re}}}^{k_{\mathrm{e}}}\frac{\mathrm{d}k}{k}\Omega_{{}_{\mathrm{GW}}}(k)h^{2}$ $\simeq c_{g}\,\Omega_{{}_{\mathrm{R}}}h^{2}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\Gamma^{2}\left(\frac{5+3w_{\mathrm{re}}}{2+6w_{\mathrm{re}}}\right)$
$\quad\times\frac{(1+3w_{\mathrm{re}})^{1+4/(1+3w_{\mathrm{re}})}}{2\pi(3w_{\mathrm{re}}-1)}\left(\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}\right)^{\frac{6w_{\mathrm{re}}-2}{1+3w_{\mathrm{re}}}},$
$\simeq c_{g}\,\Omega_{{}_{\mathrm{R}}}h^{2}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\zeta(w_{\mathrm{re}})\left(\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}\right)^{\frac{6w_{\mathrm{re}}-2}{1+3w_{\mathrm{re}}}}$ (18)
in Eq. (18), the ratio between $k_{\mathrm{e}}$ and $k_{\mathrm{re}}$ can be expressed in terms of $H_{{}_{\mathrm{I}}}$, $w_{\mathrm{re}}$ and $T_{\mathrm{re}}$ as
$\frac{k_{\mathrm{e}}}{k_{\mathrm{re}}}=\left(\frac{90}{\pi^{2}g_{\rm re}}\frac{M_{{}_{\mathrm{Pl}}}^{2}H_{{}_{\mathrm{I}}}^{2}}{T_{\mathrm{re}}^{4}}\right)^{\frac{1+3\,w_{\mathrm{re}}}{6\,(1+w_{\mathrm{re}})}}.$ (19)
@@
$\quad\times\left(\frac{c_{g}\Omega_{{}_{\mathrm{R}}}h^{2}\,\zeta(w_{\mathrm{re}})}{5.61\times 10^{-6}\,\Delta N_{\rm eff}}\frac{H_{{}_{\mathrm{I}}}^{2}}{12\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\right)^{\frac{3\,(1+w_{\mathrm{re}})}{4\,(3\,w_{\mathrm{re}}-1)}}.$ (20)
In Eq. (20), $\Delta{N}_{\mathrm{eff}}$ is tightly constrained by recent ACT data, yielding an upper limit of $\Delta{N}_{\mathrm{eff}}\leq 0.17$ at $95\%$ confidence level [8, 9]. It is important to note that there also exists a lower bound on the reheating temperature, set by BBN, which requires $T_{\mathrm{re}}\gtrsim 4$ MeV [56, 57, 58]. It can be shown that the bound on $T_{\mathrm{re}}$ derived from Eq. (20) is stronger than the BBN bound only for equations of state with $w_{\mathrm{re}}>0.58$. Therefore, in our analysis, we consider the minimum reheating temperature to be max[$T_{\mathrm{re}},T_{{}_{\mathrm{BBN}}}$], while the maximum reheating temperature corresponds to the case of instantaneous reheating, which can be directly computed from the energy scale at the end of inflation. With these tools in hand, we now turn to a discussion of the slow roll models of interest in the next section.